the vapour pressure of h20 at 20degrees Celsius is 17mmhg.calculate the vapour pressure of solution of 3g of urea in 50g of water
1
Expert's answer
2013-07-11T09:52:11-0400
To solve this we can use Raoult's law: p = p'x
x = x(H2O) = 50/18 = 2.78 moles;
moles of urea(CH4N2O) = 3/60 = 0.05 moles;
mole fraction of water in solution = 2.78/(2.78 + 0.05) = 0.98
p'x = 17*0.98 = 16.66